OpenStudy (confusedstudenthere):

Solve for x. put your answer in interval notation using grouping symbols x^2+6x<0

OpenStudy (mathstudent55):

To solve a polynomial inequality do this: 1. change the inequality sign to an equal sign and solve the equation. 2. plot each solution to the equation on a number line. 3. the number line is now divided into regions. 4. test a point from each region in the original inequality.

OpenStudy (confusedstudenthere):

You lost me at step 3

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Let's say an equation has the roots -2 and 6. You plot them as shown above.

OpenStudy (mathstudent55):

There are three regions in the number line. You can call them intervals. \((-\infty, -2)\), \((-2, 6)\), \((6, \infty)\),

OpenStudy (mathstudent55):

Now you test a point from each interval in the original inequality. If the point works in the original inequality, then that entire interval works and is part of the solution.

OpenStudy (sshayer):

\[x^2+6x<0\] add 9 to both sides of enequality \[x^2+6x+9<9\] \[\left( x+3 \right)^2<3^2\] \[\left| x+3 \right|<3\] -3<x+3<3 -6<x<0